An electric field of 1.92 kV/m and a magnetic field of 0.493 T act on a moving electron to produce no net force. If the fields are perpendicular to each other, what is the electron's speed?

Chapter 28, Problem 008 An electric field of 1.92 kV/m and a magnetic field of 0.493 T act on a moving electron to produce no net force. If the fields are perpendicular to each other, what is the electrons speed? Number Units CE the tolerance is +/-9%
Answer #1

Given that

electric field E=1.92*10^3 V/m

magnetic field B=0.493 t

basing on the concept of electro magnetic waves

now we find the electron speed

th electrons speed V=E/B=1.92*10^3/0.493=3.895*10^3 m/s